Find the values for m and b in the following systems so that the solution to the system is (1,2).
mx + 3y =8
-3x + 4y = b
Find the values for m and b in the following systems so that the solution to the systme is (-3,4).
5x + 7y = b
mx + y = 22

Answer by stanbon(75887)

Find the values for m and b in the following systems so that the solution to the system is (1,2).
mx + 3y =8
-3x + 4y = b
-----------------
I'll do one and leave the 2nd one to you.
Substitute the x=1 and y=2 into both equations
to get two equations: one in m, the other in b.
m*1+3*2 = 8
m + 6 = 8
m = 2
--------
-3*1+4*2 = b
-3+8 = b
b = 5
----------
Cheers,
Stan H.







Find the values for m and b in the following systems so that the solution to the system is (-3,4).
5x + 7y = b
mx + y = 22
Follow the same procedure to solve for m and b
You should get b = 13 and m = -6
==========
Cheers,
Stan H.